Hi plese help me wiyth this


My father is 42 and my dog is eight. If my dog were a human it would be 56. How old would my dad be if were a dog



thanks for any help coming on my way





Answer by checkley71(8403)

56/8=42/x
56x=336
x=336/56
x=6 years old for dad the dog!!!